6. Assume the length of salt is 8 bits and the dictionary for dictionary attack contain 300 words. Calculate the search space for attacking salted passwords. Give an exact number.

Ans:

7. Assume receiver B received a signed document from sender A under the digital signature on the document?

A’s public key, B’s public key, A’s private key, B’s private key

8. ________ refer to technique involving the use of human insiders to circumvent computer.

9. Encrypt “SECURITY” using Caesar cipher with k=4.

Ans:

10. which of the following attacks can harm confidentially. Choose all that apply:

Alternation, Repudiation, Masquerading, Eavesdropping, Denial- of-service

11. The output of the cryptographic hashing is called message digest.

True/False

12. Connect each of the security principles to the most relevant keyword.

___ No bypass.                                  1. Psychological acceptability.

___ Logging(auditing)                     2. Compromise recording

___User interfaces.                          3. Economy of mechanism

___Simplicity                                     4. Complete mediation
Answer #1

6.

Length of the salt = 8 bits

Total possible salts = 28 = 64

Size of dictionary = 300 words

Size of the search space = 64 x 300 = 19200

7.

The digital signature is signed using A's private key

8.

Insider Attack

9.

The given plaintext is SECURITY

k=4

The ciphertext will be

( 18 + 4 ) mod 26 = 22 = w

( 4 + 4 ) mod 26 = 8 = i

( 2 + 4 ) mod 26 = 6 = g

( 20 + 4 ) mod 26 = 24 = y

( 17 + 4 ) mod 26 = 21 = v

( 8 + 4 ) mod 26 = 12 = m

( 19 + 4 ) mod 26 = 23 = x

( 24 + 4 ) mod 26 = 2 = c

So the ciphertext is wigyvmxc

10.

Eavesdropping harms confidentiality

11.

True

12.

No bypass------------Psychological acceptability.

Logging---------------Compromise recording

User interfaces-----Economy of mechanism

Simplicity------------Complete mediation